OpenStudy (ttop0816):

Find f(-7) if f(x) = 2x + 8. -6 2 3 22

OpenStudy (anonymous):

are the numbers below possible answers?

hartnn (hartnn):

to get f(-7) just plug in x=-7 in f(x)

hartnn (hartnn):

so what u get for 2(-7) +8 =... ?

OpenStudy (anonymous):

u have to substitute (-7) in place of x. f(-7)=2(-7)+8 =-14+8 =-6